# Enhancing Cross-Domain Recommendations with Memory-Optimized LLM-Based User Agents

### v1

@@ Abstract.
Large Language Model (LLM)-based user agents have emerged as a powerful tool for improving recommender systems by simulating user interactions. However, existing methods struggle with cross-domain scenarios due to inefficient memory structures, leading to irrelevant information retention and failure to account for social influence factors such as popularity. To address these limitations, we introduce AgentCF++, a novel framework featuring a dual-layer memory architecture and a two-step fusion mechanism to filter domain-specific preferences effectively. Additionally, we propose interest groups with shared memory, allowing the model to capture the impact of popularity trends on users with similar interests. Through extensive experiments on multiple cross-domain datasets, AgentCF++ demonstrates superior performance over baseline models, highlighting its effectiveness in refining user behavior simulation for recommender systems. Our code is available at https://anonymous.4open.science/r/AgentCF-plus.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems play a pivotal role in the dissemination of information today (Ricci et al., 2010; Xia et al., 2022; Liu et al., 2023b, d, a, c, 2024, 2022). However, their development is hindered by challenges in effectively understanding user behavior (Shani and Gunawardana, 2011). One promising approach to overcoming these challenges is the reliable simulation of user interaction behavior in a controlled, privacy-preserving manner, thereby improving recommender systems by offering insights into user preferences and system performance (Zhang et al., 2024e). Recent advancements in large language models (LLMs), renowned for their capabilities in understanding, reasoning, and generating content (Zhao et al., 2023), have inspired significant efforts to develop LLM-based agents (Wang et al., 2024a). These agents often incorporate memory modules (Zhang et al., 2024b), utilize external tools (Yuan et al., 2024), and engage in advanced reasoning (Huang et al., 2024), enabling them to exhibit emergent human-like behaviors (Park et al., 2023). In this context, researchers have begun investigating the potential of LLM-based agents to simulate user interaction behavior in the field of recommender systems (Wang et al., 2024c; Zhang et al., 2024c, d; Shu et al., 2024; Cai et al., 2024; Zhang et al., 2024f).
@@
In this paper, we present AgentCF++, an enhanced version of AgentCF (Zhang et al., 2024d). Our approach introduces a dual-layer memory architecture comprising domain-separated memory and domain-fused memory, designed to prioritize target-domain-relevant information in decision-making for cross-domain scenarios. To refine this architecture, we propose an attention-inspired two-step fusion mechanism. This mechanism first identifies valuable cross-domain knowledge pertinent to the target domain and then integrates these critical preferences. Furthermore, we introduce the concept of interest groups and propose a group-shared memory mechanism to facilitate the transfer of popularity effects within the same interest group. By utilizing interest groups, we ensure that a user’s behavior impacts only those with similar interests, effectively preventing the spread of popularity effects to unrelated users. Our experimental results on five cross-domain datasets demonstrate the effectiveness of the proposed modules. Additionally, several case studies underscore the enhanced capabilities of AgentCF++.
@@ 2.2. Limitations of AgentCF
AgentCF employs a single memory for each user agent and item agent. In cross-domain scenarios, the propagation process integrates information from multiple domains into the memory of each user and item. On one hand, the mixing of cross-domain preferences in the user agent may introduce noise, thereby complicating decision-making on target domain. On the other hand, such intermingling may cause the item agent to lose its original domain characteristics. We illustrate this with a case study in Section 4.3.
@@ 3.1. Memory Architecture
Table 1. Cross-domain dataset construction.
Cross-domain Dataset Source Datasets
Cross-1 CDs & Movies & Games
Cross-2 Books & Movies & Games
Cross-3 Books & CDs & Games
Cross-4 Books & CDs & Movies
Cross-5 Books & CDs & Movies & Games
@@ 4.1.1. Datasets
We experimented with the Amazon review dataset (Hou et al., 2024a). As summarized in Table 1, we constructed cross-domain datasets Cross-1Cross-5 by combining datasets from the Books, CDs, Movies, and Games domains. Then, we retained interaction records with ratings $\geq 4$ and timestamps spanning six months, from October 2021 to March 2022. We further filtered the data to include only records of users who interacted across multiple domains and had $\geq$ 10 total interactions. Following AgentCF, we randomly sampled 100 users to minimize API call expenses. Next, we sorted these interaction records chronologically and split them into training, validation, and test sets with an 8:1:1 ratio.
@@ 4.2. Overall Performance
4.3. Case Studies
4.3.1. Memory
Figure 2 (a) and (c) illustrate the memory of AgentCF++ and AgentCF in the Game & Book domains, respectively. It is evident that the user agent in AgentCF intermixes preferences from different domains, with its memory containing both game-related preferences (e.g., role-playing) and book-related preferences (e.g., economics). As memory propagates, this mixing effect further causes its item agent to gradually lose domain-specific information.
In contrast, AgentCF++ separates preferences by domain and fuse information via a two-step fusion mechanism. First, as shown in Figure 2 (b), this mechanism extracts target-domain preferences from other domains based on cross-domain mappings (e.g., role-playing and open-world in the game domain correspond to fantasy and adventure in the book domain). Then, the mechanism fuses the extracted preferences to ensure broad coverage across domains while maintaining domain specificity. This mechanism ensures that the item agent retains information relevant to its domain.
4.3.2. Interest Group
Figure 2 (d) illustrates two interest groups along with their agenda items, each named to represent its focus.

### v2

# AgentCF++: Memory-enhanced LLM-based Agents for Popularity-aware Cross-domain Recommendations

@@ Abstract.
LLM-based user agents, which simulate user interaction behavior, are emerging as a promising approach to enhancing recommender systems. In real-world scenarios, users’ interactions often exhibit cross-domain characteristics and are influenced by others. However, the memory design in current methods causes user agents to introduce significant irrelevant information during decision-making in cross-domain scenarios and makes them unable to recognize the influence of other users’ interactions, such as popularity factors. To tackle this issue, we propose a dual-layer memory architecture combined with a two-step fusion mechanism. This design avoids irrelevant information during decision-making while ensuring effective integration of cross-domain preferences. We also introduce the concepts of interest groups and group-shared memory to better capture the influence of popularity factors on users with similar interests. Comprehensive experiments validate the effectiveness of AgentCF++. Our code is available at https://github.com/jhliu0807/AgentCF-plus.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730161isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems
@@ 1. Introduction
Recommender systems play a pivotal role in the dissemination of information today (Ricci et al., 2010; Xia et al., 2022; Liu et al., 2023b, d, a, c, 2024, 2022; Li et al., 2024). However, their development is hindered by challenges in effectively understanding user behavior (Shani and Gunawardana, 2011). One promising approach to overcoming these challenges is the reliable simulation of user interaction behavior in a controlled, privacy-preserving manner, thereby improving recommender systems by offering insights into user preferences and system performance (Zhang et al., 2024e). Recent advancements in large language models (LLMs), renowned for their capabilities in understanding, reasoning, and generating content (Zhao et al., 2023), have inspired significant efforts to develop LLM-based agents (Wang et al., 2024a). These agents often incorporate memory modules (Zhang et al., 2024b), utilize external tools (Yuan et al., 2024), and engage in advanced reasoning (Huang et al., 2024), enabling them to exhibit emergent human-like behaviors (Park et al., 2023). In this context, researchers have begun investigating the potential of LLM-based agents to simulate user interaction behavior in the field of recommender systems (Wang et al., 2024c; Zhang et al., 2024c, d; Shu et al., 2024; Cai et al., 2024; Zhang et al., 2024f).
@@
In this paper, we present AgentCF++, an enhanced version of AgentCF (Zhang et al., 2024d). Our approach introduces a dual-layer memory architecture comprising domain-separated memory and domain-fused memory, designed to prioritize target-domain-relevant information in decision-making for cross-domain scenarios. To refine this architecture, we propose an attention-inspired two-step fusion mechanism. This mechanism first identifies valuable cross-domain knowledge pertinent to the target domain and then integrates these critical preferences. Furthermore, we introduce the concept of interest groups and propose a group-shared memory mechanism to facilitate the transfer of popularity effects within the same interest group. By utilizing interest groups, we ensure that a user’s behavior impacts only those with similar interests, effectively preventing the spread of popularity effects to unrelated users. Our experimental results on five cross-domain datasets demonstrate the effectiveness of the proposed modules.
@@ 2.2. Limitations of AgentCF
AgentCF employs a single memory for each user agent and item agent. In cross-domain scenarios, the propagation process integrates information from multiple domains into the memory of each user and item. On one hand, the mixing of cross-domain preferences in the user agent may introduce noise, thereby complicating decision-making on target domain. On the other hand, such intermingling may cause the item agent to lose its original domain characteristics.
@@ 4.1.1. Datasets
We experimented with the Amazon review dataset (Hou et al., 2024a). We constructed the cross-domain datasets Cross-1Cross-5 by combining data from the Books, CDs, Movies, and Games domains, selecting 3 or 4 domains for each dataset. Then, we retained interaction records with ratings $\geq 4$ and timestamps spanning six months, from October 2021 to March 2022. We further filtered the data to include only records of users who interacted across multiple domains and had $\geq$ 10 total interactions. Following AgentCF, we randomly sampled 100 users to minimize API call expenses. Next, we sorted these interaction records chronologically and split them into training, validation, and test sets with an 8:1:1 ratio.
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China (NSFC) under the Grant No. 62172106. Peng Zhang is a faculty of School of Computer Science, Fudan University. Tun Lu is a faculty of School of Computer Science, Shanghai Key Laboratory of Data Science, Fudan Institute on Aging, MOE Laboratory for National Development and Intelligent Governance, and Shanghai Institute of Intelligent Electronics & Systems, Fudan University.